OpenStudy (anonymous):

solve system using substitution x=2y-1 2x-y=4

OpenStudy (anonymous):

We can substitute '2y - 1' for 'x' in the 2nd equation: 2x - y = 4 2(2y - 1) - y = 4 Distribute 2 into the parenthesis: 4y - 2 - y = 4 Combine like terms, can you subtract 4y - y?

OpenStudy (anonymous):

4y - 1y It's the same thing as 4 - 1, but we put a 'y' next to it. For example: 23x - 14x = 9x, because 23 - 14 = 9.

OpenStudy (anonymous):

3y?

OpenStudy (anonymous):

Yes! So we have: \(3y - 2 = 4\) Now add 2 to both sides, what's 4 + 2?

OpenStudy (anonymous):

3y=6?

OpenStudy (anonymous):

Yep, now divide 3 to both sides.

OpenStudy (anonymous):

y=3?

OpenStudy (anonymous):

Yes, so we have the y-value of the solution, now we can plug this into any of the two equations to find the x-value. x = 2y - 1 x = 2(3) - 1 Solve for 'x'.

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

Yes, so our full solution is: (5, 3)

OpenStudy (anonymous):

x=2y-1 then x=2(2x-4)-1 then x=4x-8-1 x=4x-9 then x-4x=-9 then -3x=-9 then x=3 then solve for y

OpenStudy (anonymous):

Welcome to Open Study! You can give medals by clicking 'Best Resposne'. @priea

OpenStudy (anonymous):

thanks but I put that in and it said it was wrong. can u help me with another one please

OpenStudy (anonymous):

Sure, close this one and open a new one.

OpenStudy (anonymous):

Oh, you divided wrong.. 6 / 3 isn't 3..it's 2. So the 'y' would actually be '2'. Then we plug that in and get '3' for 'x'. So our answer would be (3, 2)..